40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 748     Entry time: Wed Dec 7 20:44:36 2011
Author: Frank 
Type: DailyProgress 
Category: RFAM 
Subject: temporary channels 

i'm currently adding new channels for RFAM monitoring and EOM temp control but before they work i used some other channels for now.

  • The EOM temp setpoint is currently modulated using the SR785 as i cant add a large DC offset on a small modulation using the SR345.
  • The current setpoint is 650mV which is equivalent to something around 30C.
  • On top i have a modulation (sine, 1mHz) of 180mV (?, have to double check, might be 170mV) to modulate the EOM temperature.
  • The setpoint value is recorded using channel C3:PSL-FSS_VCOMON
  • The actual Temperature of the in-loop sensor is recorded using C3:PSL-FSS_FREQCOUNT
  • The feedback to the laser is monitored using the fast actuator channel C3:PSL-FSS_FAST. The EOM feedback loop is disabled as well as the script for laser temp feedback.

i checked the temp modulation by using a PT1000 sensor as an OOL sensor on the EOM case using a multimeter.with min/max hold function.

  • min value is 1120 Ohms = 30.9C
  • max value is 1139 Ohms = 35.8C

Monitoring the the above channels using Striptool didn't show any result yet. Even the modulation of ~5K is not visible in the feedback to the laser. Will add all signals to their final channels tomorrow and do a more detailed analysis.

ELOG V3.1.3-